okay, here is another one. I am not understanding the whole exponent thing.
Simplify each of the following expressions. Write your answers with positive exponents only.
(m^4)^-3/(m^-2)^4
The book shows two different examples according to different properties, how do I know which one to apply?? I am really confused with the properties overall.
Help!

Answer by checkley75(3666)

I MUST APPOLOGIZE FOR INCORRECT DATA. iT WAS LATE & I WASN'T THINKING STRAIGHT.
WHEN YOU RAISE POWERS BY POWERS YOU NEED TO MULTIPLY THE POWERS.
THUS (M^4)^-3=M^(4*-3)=M^-12
(M^-2)^4=M^(-2*4)=M^-8
NOW WE HAVE:
M^-12/M^-8
1/M^12*M^8
M^8/M^12
M^(8-12)=M^-4 OR 1/M^4